# High-Order Hermite Optimization: Fast and Exact Gradient Computation in Open-Loop Quantum Optimal Control using a Discrete Adjoint Approach

### v2

@@ 4.1 Correctness Experiments on a Rabi Oscillator Problem
To verify the correctness of our numerical method, we simulate a Rabi oscillator problem, which has an analytic solution. A Rabi oscillator problem consists of a two-level system in a rotating frame of reference. For this problem, Schrödinger’s equation is
@@
$U(t)=\begin{bmatrix}\cos(|\Omega|t)&(\sin(\theta)-i\cos(\theta))\sin(|\Omega|t)\\ -(\sin(\theta)+i\cos(\theta))\sin(|\Omega|t)&\cos(|\Omega|t)\end{bmatrix},\quad\text{where }\Omega=|\Omega|(\cos(\theta)+i\sin(\theta)).$ (35)
This solution can be derived using an eigenvector decomposition of the Hamiltonian [31, Sec. 7.23]. The populations of the $|0\rangle$ and $|1\rangle$ states oscillate between 0 and 1 with period $\tau=\pi/|\Omega|$, and hence an application of control pulses $p_{0}={\rm Re}\,(\Omega)$ and $q_{0}={\rm Im}\,(\Omega)$ for duration $\tau$ is called a Rabi oscillation. A control pulse with $p_{0}=|\Omega|$, $q_{0}=0$, and duration $\tau/2$ implements a Pauli-$X$ gate.
@@ 4.1.1 Verification of Order of Accuracy
Using the analytic solution (35), we may evaluate the error in a numerical solution of the Rabi oscillator problem (34) at any point in time for any initial condition. In Table 2, we plot the Frobenius norm of the relative error in the final-time numerical simulation of the Rabi oscillator problem for a basis of initial states $U_{0}=\operatorname{diag}(1,1)$, with $\Omega=0.05(\cos(\pi/4)+i\sin(\pi/4))$ (constant control pulses with amplitude $25/\sqrt{2}$ MHz), and $T=9.5\pi/|\Omega|$ (nine and a half Rabi oscillations), starting with sixteen timesteps and then doubling the number of timesteps until we reach 256 timesteps. We observe that the method converges to the analytic solution with the expected orders of accuracy. Furthermore, we can simulate the problem with high accuracy using far fewer timesteps with the high-order methods than with the second-order method.
@@
Table 2: Frobenius norms and convergence rates of the final-time relative error in the numerical solution of nine and a half periods of Rabi oscillation, using method orders two through twelve.
@@ 4.1.2 Verification of Exact Gradients
We can use the Rabi oscillator in an optimal control problem by taking $\boldsymbol{\theta}=[p_{0},q_{0}]^{T}$ and $\Omega=p_{0}+iq_{0}$ so that $H_{c}=p_{0}(a+a^{\dagger})+q_{0}(a-a^{\dagger})$, and assigning some target gate $U^{\textrm{target}}\in\mathbb{C}^{2\times 2}$. We can calculate the derivatives of each entry of $U(t)$ with respect to $p_{0}$ and $q_{0}$ by applying straightforward calculus to (35):
$\frac{\partial U(t)}{\partial p_{0}}=\begin{bmatrix}-\frac{p_{0}t}{|\Omega|}\sin(|\Omega|t)&\frac{-1}{p_{0}-iq_{0}}\left(\frac{q_{0}}{|\Omega|}\sin(|\Omega|t)+ip_{0}t\cos(|\Omega|t)\right)\\ \frac{1}{p_{0}+iq_{0}}\left(\frac{q_{0}}{|\Omega|}\sin(|\Omega|t)-ip_{0}t\cos(|\Omega|t)\right)&-\frac{p_{0}t}{|\Omega|}\sin(|\Omega|t)\end{bmatrix},$ (36)
$\frac{\partial U(t)}{\partial q_{0}}=\begin{bmatrix}-\frac{q_{0}t}{|\Omega|}\sin(|\Omega|t)&\frac{1}{q_{0}+ip_{0}}\left(q_{0}t\cos(|\Omega|t)+\frac{ip_{0}}{|\Omega|}\sin(|\Omega|t)\right)\\ \frac{-1}{p_{0}+iq_{0}}\left(\frac{p_{0}}{|\Omega|}\sin(|\Omega|t)+iq_{0}t\cos(|\Omega|t)\right)&-\frac{q_{0}t}{|\Omega|}\sin(|\Omega|t)\end{bmatrix}.$ (37)
@@
$\frac{d\mathcal{I}}{d{p_{0}/q_{0}}}=-\frac{1}{2}\bigg{(}\left\langle U^{\textrm{target}}_{\rm Re}\,,U_{T,{\rm Re}\,}\right\rangle_{F}+\left\langle U^{\textrm{target}}_{\rm Im}\,,U_{T,{\rm Im}\,}\right\rangle_{F}\bigg{)}\left(\left\langle U^{\textrm{target}}_{\rm Re}\,,\frac{dU_{T,{\rm Re}\,}}{d{p_{0}/q_{0}}}\right\rangle_{F}+\left\langle U^{\textrm{target}}_{\rm Im}\,,\frac{dU_{T,{\rm Im}\,}}{d{p_{0}/q_{0}}}\right\rangle_{F}\right)\\ -\frac{1}{2}\bigg{(}\left\langle U^{\textrm{target}}_{\rm Re}\,,U_{T,{\rm Im}\,}\right\rangle_{F}-\left\langle U^{\textrm{target}}_{\rm Im}\,,U_{T,{\rm Re}\,}\right\rangle_{F}\bigg{)}\left(\left\langle U^{\textrm{target}}_{\rm Re}\,,\frac{dU_{T,{\rm Im}\,}}{d{p_{0}/q_{0}}}\right\rangle_{F}-\left\langle U^{\textrm{target}}_{\rm Im}\,,\frac{dU_{T,{\rm Re}\,}}{d{p_{0}/q_{0}}}\right\rangle_{F}\right),$
where the sensitivites ${dU_{T,{\rm Re}\,/{\rm Im}\,}}/{d{p_{0}}}$ and ${dU_{T,{\rm Re}\,/{\rm Im}\,}}/{d{q_{0}}}$ may be computed explicitly using (36) and (37).
To test the correctness of our discrete adjoint implementation, we take the Hadamard gate as our target gate and again use $\Omega=p_{0}+iq_{0}=0.05(\cos(\pi/4)+i\sin(\pi/4))$ and $T=9.5\pi/|\Omega|$. We compute the gradient using the discrete adjoint method for 16 to 256 timesteps and method order 2 through 12, and compare the Frobenius norm of the difference between the analytic gradient of the continuous problem and the gradient of the discrete problem using the discrete adjoint method. The results are shown in Table 3. We observe that as the stepsize increases, the gradient converges to the analytic gradient with machine precision.
@@
Table 3: $\ell_{2}$ norms and convergence rates of the error in the discrete-adjoint-computed gradient of the infidelity for the Rabi oscillator problem with the Hadamard gate as the target. Numerical solution done over nine and a half periods of Rabi oscillation, using method orders two through twelve.
@@ 4.2 Speedup Experiments on a Multi-Qudit Gate Design Problem
As a practical example, we consider a two qudits coupled to a resonator bus. Using the model (4), the drift Hamiltonian is
@@
In the model, the resonator is treated no differently than another qudit, but we use the index $R$ to indicate its significance as a resonator: the two qudits used for computation are weakly coupled to one another, but more strongly coupled to the resonator, which allows the resonator to mediate the interaction between the two qudits. For the physical parameters, we use the values from the CNOT3 example of Juqbox.jl [2], which we provide in Table 4. We also mimic Juqbox.jl’s strategy of weighting the guard penalty for the guard states exponentially based on the distance from the highest level guard state. Specifically, the weight assigned to a guard state is $0.001^{n_{G}}/N_{G}$, where $n_{G}$ is the minimum number of levels difference between the the current state and the highest guard state in each subsystem (e.g. for the tripartite system, for the state $|100\rangle$ we get $n_{G}=8$, and for $|121\rangle$ we get $n_{G}=0$), and $N_{G}$ is the total number of guard levels in the computational basis (i.e., the dimension of the guard subspace), or $N_{G}=79$ for the tripartite system. This allows optimization to populate the lower levels of the cavity, which may which advantageous for facilitating the interaction of the two qudits, while more heavily discouraging the population of the guard states which push the limits of the model using a truncated Hilbert space.
@@
Value$/2\pi$ (GHz) 4.11 4.82 7.84 1.20(-2) 2.25(-1) 2.83(-5) 1.00(-6) 2.49(-3) 2.52(-3)
Table 4: Physical parameters used in (38) to model the Hamiltonian of two qubits coupled to a resonator bus, rounded to three significant digits. The qubits are weakly coupled to one another, but more strongly coupled to the resonator.
@@
We model each of the two qubits as four-level subsystems, and the resonator as a ten-level subsystem. We wish to implement a CNOT gate on the two qudits, utilizing the resonator to mediate the interaction. We use two essential levels and two guard levels to the qudits, and one essential level and nine guard levels to the resonator (see Table 5).
Qudit $1$ Qudit $2$ Resonator
# Essential Levels 2 2 1
# Guard Levels 2 2 9
Table 5: Number of essential and guard levels for each subsystem in model of two qudits coupled to a resonator bus.
@@
We fix the duration of the gate to 550 nanoseconds. As our control pulse ansatz, we use carrier waves multiplying degree 14 B-spline curves with 15 B-spline wavelets for the real and imaginary parts of the envelope, for a total of $3\times 3\times 2\times 15=270$ control parameters. For each control Hamiltonian corresponding to a subsystem, we use three carrier wave frequencies: 00, $-2\pi\xi_{1}$, and $-2\pi\xi_{2}$ for qudits $1$ and $2$, and 00, $-2\pi\xi_{R1}$, and $-2\pi\xi_{R2}$ for the resonator. The B-spline degree of 14 ensures that the Hamiltonian is smooth enough that the methods up to order 12 will exhibit their full order of accuracy in the asymptotic region. For the lower-order methods, we could lower the degree of the B-spline curves to reduce the cost of evaluating the control functions while keeping the full order of accuracy, but we use the same degree for all orders for the sake of keeping the comparison as direct as possible.
@@ 4.2.1 Determining Appropriate Step Size
In order for the optimization to find control pulses which implement a gate with low infidelity, the numerical solution of Schrödinger’s equation must be accurate enough to ensure that the error in the infidelity is reasonably smaller than the infidelity itself. With this in mind, before performing the optimization using the Hermite method of order $N$ we perform a numerical experiment to determine an appropriate stepsize $\Delta t$ to use in order to reach a target relative error in the numerical solution of Schrödinger’s at the final time of the gate.
For each order of the method, starting with two timesteps and doubling the number of timesteps each iteration until the state history cannot fit into main memory, we perform a forward evolution, record the final state, perform the rest of the gradient computation (adjoint evolution and gradient accumulation), and record the time taken for the full gradient computation. We do this for 25 samples of the control vector $\boldsymbol{\theta}$ from a uniform distribution with a range of $[-0.05,0.05]$ for each control parameter, which enforces a maximum amplitude of 50 MHz on the B-spline envelope for each carrier wave for each control pulse. For each order of the method, the same 25 samples are used, so that each method is attempting to simulate the exact same analytic dynamics. An absolute tolerance of $10^{-15}$ was for the linear solves, which were performed using GMRES. This tolerance could be relaxed for greater computational efficiency, depending on the accuracy required of the state vector. In Figure 3, we report the number of timesteps, mean time to compute the gradient, and mean relative error at the final time for each order of the method. The samples were collected using a single core of a Intel® Xeon® Gold 6148 CPU clocked at 2.40GHz, and 32 GB of memory, in Michigan State University’s high-performance computing cluster.
@@
In order to estimate the number of timesteps needed to reach a target relative error, we use linear interpolation of the number of timesteps and the mean relative error in the case that we have recorded mean relative errors above and below the target relative error. Otherwise, we estimate the number of timesteps using a least squares fit of the the number of timesteps and the mean relative error in the region of asymptotic convergence.
This heuristic approach warrants some skepticism that we can use a single stepsize in order to consistently achieve a relative error below the target relative error during the optimization, since the control vector (and hence the dynamics) changes with each iteration of the optimization. However, in our experiment we observe that the standard deviation of the relative error is always an order of magnitude smaller than the mean relative error, indicating that the magnitude of the relative error is consistent for different values of $\boldsymbol{\theta}$. We use this numerical observation to justify using a single stepsize during the optimization. Table 6 shows the number of timesteps required for to reach several target relative final-time errors for each order of the Hermite method, and Table 7 shows the expected speedup of the gradient computation (to perform forward evolution, adjoint evolution, and gradient accumulation) for each order of the HOHO method compared to Störmer-Verlet, as implemented by Juqbox.jl, based on the average time taken to compute the gradient over the 25 different control vectors.
@@
Table 6: The number of timesteps for each order of the Hermite method (and for Störmer-Verlet) required to reach a target relative final-time error in the numerical solution of Schrödinger’s equation for the three subsystem example, starting in the initial states of the gate design problem: $|000\rangle,|001\rangle,|010\rangle$, and $|011\rangle$. Results are based on the average accuracy for 25 randomly sampled control vectors.
@@
Table 7: The speedup factor for each order of the Hermite method over Störmer-Verlet for the three subsystem example, starting in the initial states of the gate design problem: $|000\rangle,|001\rangle,|010\rangle$, and $|011\rangle$. Results are based on the average time elapsed for 25 randomly sampled control vectors.
In Tables 8 and 9, we again show the number of timesteps required and the average speedup for the three subsystem example, but for the evolution of a single initial state: $|933\rangle$, the most highly excited state of the system, for which the dynamics are expected to happen over much shorter timescales. This represents a worst-case scenario computationally, as more timesteps are required to resolve the faster dynamics, and consequently the higher-order methods are much more efficient, since their accuracy more significantly with the number of timesteps. Fast dynamics could also occur even in the computational subspace when there are many qudits, for example, or if the computation is done in the laboratory frame instead of the rotating frame.
@@
Table 8: The number of timesteps for each order of the Hermite method (and for Störmer-Verlet) required to reach a target relative final-time error in the numerical solution of Schrödinger’s equation for the three subsystem example, starting in the most highly excited state: $|933\rangle$. Results are based on the average accuracy for 25 randomly sampled control vectors.
@@
Table 9: The speedup factor for each order of the Hermite method over Störmer-Verlet for the three subsystem example, starting in the most highly excited state: $|933\rangle$. Results are based on the average time elapsed for 25 randomly sampled control vectors.

### v3

@@ 4.1 Correctness Experiments on a Rabi Oscillator Problem
To verify the correctness of our numerical method, we simulate a Rabi oscillator problem, which has an analytical solution. A Rabi oscillator problem consists of a two-level system in a rotating frame of reference. For this problem, Schrödinger’s equation is
@@
$U(t)=\begin{bmatrix}\cos(|\Omega|t)&(\sin(\varphi)-i\cos(\varphi))\sin(|\Omega|t)\\ -(\sin(\varphi)+i\cos(\varphi))\sin(|\Omega|t)&\cos(|\Omega|t)\end{bmatrix},\quad\text{where }\Omega=|\Omega|(\cos(\varphi)+i\sin(\varphi)).$ (35)
This solution can be derived using an eigenvector decomposition of the Hamiltonian [40, Sec. 7.23]. The populations of the $|0\rangle$ and $|1\rangle$ states oscillate between 0 and 1 with period $\tau=\pi/|\Omega|$, a phenomenon known as Rabi oscillation.
@@ 4.1.1 Verification of Order of Accuracy
Using the analytical solution (35), we may evaluate the error in a numerical solution of the Rabi oscillator problem (34) at any point in time for any initial condition. In Table 3, we plot the Frobenius norm of the relative error in the final-time numerical solution of the Rabi oscillator problem for a basis of initial states $U_{0}=\operatorname{diag}(1,1)$, with $\Omega=0.05(\cos(\pi/4)+i\sin(\pi/4))$ (constant control pulses with amplitude $25/\sqrt{2}$ MHz), and $T=9.5\pi/|\Omega|$ (nine and a half Rabi oscillations), starting with sixteen timesteps and then doubling the number of timesteps until we reach 256 timesteps. We observe that the method converges to the analytical solution with the expected orders of accuracy. Furthermore, we can simulate the problem with high accuracy using far fewer timesteps with the high-order methods than with the second-order method.
@@
Table 3: Frobenius norms and convergence rates of the final-time relative error in the numerical solution of nine and a half periods of Rabi oscillation, using the Hermite method with orders two through twelve.
@@ 4.1.2 Verification of Exact Gradients
We can use the Rabi oscillator in an optimal control problem by taking $\boldsymbol{\theta}=[{\theta_{1},\theta_{2}}]^{T}$ and $\Omega={\theta_{1}+i\theta_{2}}$, so that $H_{c}={\theta_{1}}(a+a^{\dagger})+{\theta_{2}}(a-a^{\dagger})$, and assigning some target gate $U^{\textrm{target}}\in\mathbb{C}^{2\times 2}$. For example, taking $\theta_{1}=|\Omega|,$ $\theta_{2}=0$, and using a gate duration of $\tau/2=\pi/(2|\Omega|)$ implements a Pauli-$X$ gate.
We can calculate the derivatives of each entry of $U(t)$ with respect to $\theta_{1}$ and $\theta_{2}$ by applying straightforward calculus to (35):
$\frac{\partial U(t)}{\partial\theta_{1}}=\begin{bmatrix}-\frac{\theta_{1}t}{|\Omega|}\sin(|\Omega|t)&\frac{-1}{\theta_{1}-i\theta_{2}}\left(\frac{\theta_{2}}{|\Omega|}\sin(|\Omega|t)+i\theta_{1}t\cos(|\Omega|t)\right)\\ \frac{1}{\theta_{1}+i\theta_{2}}\left(\frac{\theta_{2}}{|\Omega|}\sin(|\Omega|t)-i\theta_{1}t\cos(|\Omega|t)\right)&-\frac{\theta_{1}t}{|\Omega|}\sin(|\Omega|t)\end{bmatrix}$, (36)
$\frac{\partial U(t)}{\partial\theta_{2}}=\begin{bmatrix}-\frac{\theta_{2}t}{|\Omega|}\sin(|\Omega|t)&\frac{1}{\theta_{2}+i\theta_{1}}\left(\theta_{2}t\cos(|\Omega|t)+\frac{i\theta_{1}}{|\Omega|}\sin(|\Omega|t)\right)\\ \frac{-1}{\theta_{1}+i\theta_{2}}\left(\frac{\theta_{1}}{|\Omega|}\sin(|\Omega|t)+i\theta_{2}t\cos(|\Omega|t)\right)&-\frac{\theta_{2}t}{|\Omega|}\sin(|\Omega|t)\end{bmatrix}$. (37)
@@
$\frac{d\mathcal{I}}{d{\theta_{1}/\theta_{2}}}=-\frac{1}{2}\bigg(\left\langle U^{\textrm{target}}_{\rm Re}\,,U_{T,{\rm Re}\,}\right\rangle_{F}+\left\langle U^{\textrm{target}}_{\rm Im}\,,U_{T,{\rm Im}\,}\right\rangle_{F}\bigg)\left(\left\langle U^{\textrm{target}}_{\rm Re}\,,\frac{dU_{T,{\rm Re}\,}}{d{\theta_{1}/\theta_{2}}}\right\rangle_{F}+\left\langle U^{\textrm{target}}_{\rm Im}\,,\frac{dU_{T,{\rm Im}\,}}{d{\theta_{1}/\theta_{2}}}\right\rangle_{F}\right)\\ -\frac{1}{2}\bigg(\left\langle U^{\textrm{target}}_{\rm Re}\,,U_{T,{\rm Im}\,}\right\rangle_{F}-\left\langle U^{\textrm{target}}_{\rm Im}\,,U_{T,{\rm Re}\,}\right\rangle_{F}\bigg)\left(\left\langle U^{\textrm{target}}_{\rm Re}\,,\frac{dU_{T,{\rm Im}\,}}{d{\theta_{1}/\theta_{2}}}\right\rangle_{F}-\left\langle U^{\textrm{target}}_{\rm Im}\,,\frac{dU_{T,{\rm Re}\,}}{d{\theta_{1}/\theta_{2}}}\right\rangle_{F}\right),$
where the sensitivities ${dU_{T,{\rm Re}\,/{\rm Im}\,}}/{d{\theta_{1}}}$ and ${dU_{T,{\rm Re}\,/{\rm Im}\,}}/{d{\theta_{2}}}$ may be computed explicitly using (36) and (37).
To test the correctness of our discrete adjoint implementation, we take the Hadamard gate as our target gate and again use $\Omega=\theta_{1}+i\theta_{2}=0.05(\cos(\pi/4)+i\sin(\pi/4))$ and $T=9.5\pi/|\Omega|$. We compute the gradient using the discrete adjoint method for 16 to 256 timesteps and method order 2 through 12, and compare the Frobenius norm of the difference between the analytic gradient of the continuous problem and the gradient of the discrete problem using the discrete adjoint method. The results are shown in Table 4. We observe that as the number of timesteps increases, the gradient converges to the analytic gradient with machine precision.
@@
Table 4: $\ell_{2}$ norms and convergence rates of the error in the discrete-adjoint-computed gradient of the infidelity for the Rabi oscillator problem with the Hadamard gate as the target. Numerical solution done over nine and a half periods of Rabi oscillation, using the Hermite method with orders two through twelve.
@@ 4.2 Speedup Experiments on a Multi-Qudit Gate Design Problem
As a practical example, we consider two qudits coupled to a resonator bus. Using the model (4), the drift Hamiltonian is
@@
In the model, the resonator is treated no differently than another qudit, but we use the index $R$ to indicate its significance as a resonator: the two qudits used for computation are weakly coupled to one another, but more strongly coupled to the resonator, which allows the resonator to mediate the interaction between the two qudits. For the physical parameters, we use the values from the CNOT3 example of Juqbox.jl [2], which we provide in Table 5(a).
@@
[…]
((a)) Physical parameters, rounded to three significant digits. The qubits are weakly coupled to one another, but more strongly coupled to the resonator.
Resonator Qudit 11 Qudit 22
# Essential Levels 1 2 2
# Guard Levels 9 2 2
((b)) Number of essential and guard levels for each subsystem.
Table 5: Parameters used to model the Hamiltonian of two qudits coupled to a resonator bus in (38).
@@
We model each of the two qubits as four-level subsystems (making them qudits), and the resonator as a ten-level subsystem. We wish to implement a CNOT gate on the two qudits, utilizing the resonator to mediate the interaction. We use two essential levels and two guard levels for the qudits, and one essential level and nine guard levels for the resonator (see Table 5(b)). We also mimic Juqbox.jl’s strategy of weighting the guard penalty for the guard states exponentially based on the distance from the highest level guard state. Let $N_{G}$ be the total number of guard levels in the computational basis (i.e., the dimension of the guard subspace), which is $N_{G}=156$ for this example. Let $N_{E,K}$ and $N_{G,K}$ be the number of essential and guard states in subsystem $K$ for $K\in\{1,2,R\}$. Then the weight of the guard penalty for the state $|i_{R}i_{1}i_{2}\rangle$ is $w_{G}=\max\{w_{G,1},w_{G,2},w_{G,R}\}/N_{G}$, where $w_{G,K}$ are weights associated with the leakage into the guard subspace for each subsystem. They are defined by:
$w_{G,K}=\begin{cases}0&\textrm{if }i_{K}<N_{E,K},\\ 0.001^{N_{E,K}+N_{G,K}-i_{K}-1}&\textrm{otherwise.}\end{cases}$
For example, for the state $|501\rangle$ we get $w_{G,R}=0.001^{4}$, $w_{G,1}=0$, and $w_{G,2}=0$, and therefore the weight of the guard penalty is $w_{G}=0.001^{4}/156$. For the state $|713\rangle$ we get $w_{G,R}=0.001^{2},$ $w_{G,1}=0$, and $w_{G,2}=1$, and therefore the weight of the guard penalty is $w_{G}=1/156$.
Weighting the guard levels in this way allows the optimization to populate the lower levels of the cavity, which may be advantageous for facilitating the interaction of the two qudits, while more heavily discouraging the population of the guard states which push the limits of the truncated Hilbert space used in the model.
@@
We fix the duration of the gate to 550 nanoseconds. As our control pulse ansatz, we use carrier waves multiplying degree 14 B-spline curves with 15 B-spline wavelets for the real and imaginary parts of the envelope, for a total of $3\times 3\times 2\times 15=270$ control parameters. For each control Hamiltonian corresponding to a subsystem, we use three carrier wave frequencies: 0, $2\pi\xi_{1}$, and $2\pi\xi_{2}$ for qudits $1$ and $2$, and 0, $2\pi\xi_{R1}$, and $2\pi\xi_{R2}$ for the resonator. The B-spline degree of 14 ensures that the Hamiltonian is smooth enough that the methods up to order 12 will exhibit their full order of accuracy in the asymptotic region. For the lower-order methods, we could lower the degree of the B-spline curves to reduce the cost of evaluating the control functions while keeping the full order of accuracy, but we use the same degree for all orders for the sake of keeping the comparison as direct as possible.
@@ 4.2.1 Determining Appropriate Step Size
In order for the optimization to find control pulses which implement a gate with low infidelity, the numerical solution of Schrödinger’s equation must be accurate enough to ensure that the error in the infidelity is reasonably smaller than the infidelity itself. With this in mind, before performing the optimization using the Hermite method of a given order, we perform a numerical experiment to determine an appropriate stepsize $\Delta t$ to use in order to reach a target relative error in the numerical solution of Schrödinger’s at the final time of the gate.
For each order of the method, starting with two timesteps and doubling the number of timesteps each iteration until the state history cannot fit into main memory, we perform a forward evolution, record the final state, perform the rest of the gradient computation (adjoint evolution and gradient accumulation), and record the time taken for the full gradient computation. We do this for 25 samples of the control vector $\boldsymbol{\theta}$ from a uniform distribution with a range of [0.05,0.05][-0.05,0.05] for each control parameter, which enforces a maximum amplitude of 50 MHz on the B-spline envelope for each carrier wave for each control pulse. For each order of the method, the same 25 samples are used, so that each method is attempting to simulate the exact same analytic dynamics. An absolute tolerance of $10^{-15}$ was used for the linear solves, which were performed using GMRES. This tolerance could be relaxed for greater computational efficiency, depending on the accuracy required of the state vector. In Figure 3, we report the number of timesteps, mean time to compute the gradient, and mean relative error at the final time for each order of the method. The samples were collected using a single core of an Intel® Xeon® Gold 6148 CPU clocked at 2.40 GHz, and 32 GB of memory, in Michigan State University’s high-performance computing cluster.
@@
In order to estimate the number of timesteps needed to reach a target relative error, we use linear interpolation of the number of timesteps and the mean relative error in the case that we have recorded mean relative errors above and below the target relative error. Otherwise, we estimate the number of timesteps using a least squares fit of the number of timesteps and the mean relative error in the region of asymptotic convergence.
This heuristic approach warrants some skepticism that we can use a single stepsize in order to consistently achieve a relative error below the target relative error during the optimization, since the control vector (and hence the dynamics) changes with each iteration of the optimization. However, in our experiment we observe that the standard deviation of the relative error is always an order of magnitude smaller than the mean relative error, indicating that the magnitude of the relative error is consistent for different values of $\boldsymbol{\theta}$. We use this numerical observation to justify using a single stepsize during the optimization. Table 6(a) shows the number of timesteps required for to reach several target relative final-time errors for each order of the Hermite method, and Table 6(b) shows the expected speedup of the gradient computation (to perform forward evolution, adjoint evolution, and gradient accumulation) for each order of the HOHO method compared to Störmer-Verlet, as implemented by Juqbox.jl, based on the average time taken to compute the gradient over the 25 different control vectors.
@@
((a)) Number of timesteps required to reach a target relative final-time error.
@@
((b)) Speedup factor (relative to Störmer-Verlet) for each order of the Hermite method.
Table 6: The number of timesteps required and speedup factors observed when using Störmer-Verlet and each order of the Hermite method to numerically solve Schrödinger’s equation to a target relative final-time error for the three subsystem example, starting in the initial states of the gate design problem: $|000\rangle,|001\rangle,|010\rangle$, and $|011\rangle$. Results are based on the average accuracy and elapsed time for 25 randomly sampled control vectors.
In Tables 7(a) and 7(b), we again show the number of timesteps required and the average speedup for the three subsystem example, but for the evolution of a single initial state: $|933\rangle$, the most highly excited state of the system, for which the dynamics are expected to happen over much shorter timescales. This represents a worst-case scenario computationally, as more timesteps are required to resolve the faster dynamics, and consequently the higher-order methods are much more efficient, since their accuracy increases more significantly with the number of timesteps. Fast dynamics could also occur even in the computational subspace when there are many qudits, for example, or if the computation is done in the laboratory frame instead of the rotating frame.
@@
((a)) Number of timesteps required to reach a target relative final-time error.
@@
((b)) Speedup factor (relative to Störmer-Verlet) for each order of the Hermite method.
Table 7: The number of timesteps required and speedup factors observed when using Störmer-Verlet and each order of the Hermite method to numerically solve Schrödinger’s equation to a target relative final-time error for the three subsystem example, starting in the most highly excited state: $|933\rangle$. Results are based on the average accuracy and elapsed time for 25 randomly sampled control vectors.